# Good Seed Makes a Good Crop: Discovering Secret Seeds in Text-to-Image Diffusion Models

### v1

@@ Abstract
Recent advances in text-to-image (T2I) diffusion models have facilitated creative and photorealistic image synthesis. By varying the random seeds, we can generate various images for a fixed text prompt. Technically, the seed controls the initial noise and, in multi-step diffusion inference, the noise used for reparameterization at intermediate timesteps in the reverse diffusion process. However, the specific impact of the random seed on the generated images remains relatively unexplored. In this work, we conduct a large-scale scientific study into the impact of random seeds during diffusion inference. Remarkably, we reveal that the best ‘golden’ seed achieved an impressive FID of 21.60, compared to the worst ‘inferior’ seed’s FID of 31.97. Additionally, a classifier can predict the seed number used to generate an image with over 99.9% accuracy in just a few epochs, establishing that seeds are highly distinguishable based on generated images. Encouraged by these findings, we examined the influence of seeds on interpretable visual dimensions. We find that certain seeds consistently produce grayscale images, prominent sky regions, or image borders. Seeds also affect image composition, including object location, size, and depth. Moreover, by leveraging these ‘golden’ seeds, we demonstrate improved image generation such as high-fidelity inference and diversified sampling. Our investigation extends to inpainting tasks, where we uncover some seeds that tend to insert unwanted text artifacts. Overall, our extensive analyses highlight the importance of selecting good seeds and offer practical utility for image generation.
@@ 1 Introduction
Text-to-Image (T2I) diffusion models [2, 3, 6, 31, 35, 36, 56] have advanced image synthesis significantly, enabling the creation of photorealistic, high-resolution images. However, their training requires substantial computational resources, limiting such research to a few well-equipped labs. Despite these limitations, many studies have enhanced image generation during inference by feature re-weighting [43], gradient-based guidance [10, 42, 48], or integration with multimodal LLMs [5, 54].
@@
In this work, we propose an inference technique to enhance image generation by exploring ‘secret seeds’ in the reverse diffusion process. Inspired by research like Torch.manual_seed(3407) [32], which revealed that well-chosen neural network initialization seeds can outperform poorly chosen ones in image classification, we investigate whether ‘golden’ or ‘inferior’ seeds similarly impact image quality in T2I diffusion inference. Surprisingly, using the pretrained popular T2I model Stable Diffusion (SD) 2.0 [36] across 1,024 seeds, we discovered that the best ‘golden’ seed achieved an FID [15, 40] of 21.60, whereas the worst ‘inferior’ seed only reached an FID of 31.97—a significant difference within the community. This finding sparked our curiosity to understand several scientific questions: What does the seed control in T2I diffusion inference? Why are random seeds so impactful? Can seeds be distinguished by the images they generate? Do they control interpretable image dimensions, and if so, how can this be leveraged to enhance image generation?
@@
To address our research questions, we first examined how random seeds control the initial noisy latent and the Gaussian noise during the reparameterization step of each intermediate timestep in the reverse latent diffusion process, as detailed in Section 3.1. We also developed a dataset using two T2I diffusion models: the conventional multi-step SD 2.0 [36] and the distilled one-step SDXL Turbo [38]. This dataset includes over 22,000 diverse text prompts and, using 1,024 unique fixed seeds for each combination of model and prompt, resulted in approximately 46 million images as discussed in Section 3.2. Our initial objective was to determine whether each random seed encodes unique characteristics identifiable in the generated images. To test this, we trained a 1,024-way classifier to predict the seed number used during diffusion inference from the generated images across diverse prompts. Remarkably, this classifier reached over 99.9% validation accuracy after just six epochs, a stark contrast to the random guessing chance of approximately 0.01%, establishing that seeds are highly distinguishable based on the generated images as shown in Section 3.3.
@@
Having confirmed seed distinguishability, we aim to understand if there are any interpretable perceptual dimensions enabling this differentiation. Our next step involves designing a pipeline to extract style and layout representations, apply dimensionality reduction [1, 49] for visible clustering, and identify consistent patterns across different seeds, regardless of the input prompts. For example, certain seeds consistently produce ‘grayscale’ images, others generate images with prominent white ‘sky’ regions at the top, and some seeds create image borders or insert ‘text’ during inpainting mode. In terms of image layout, various seeds consistently influence the main subject’s scale, location, and depth within the image. The details on these findings are provided in Section 3.4.
@@
Building on these discoveries from our seed analysis, we propose several downstream applications to enhance image generation, as detailed in Section 4. First, by identifying ‘golden’ seeds across a variety of prompts, we can limit sampling to the top-K seeds for high-fidelity inference. This approach demonstrates superior quantitative performance, as measured by FID [40] and HPS v2 [52], compared to random sampling in the default implementation. Second, our findings indicate that certain seeds capture distinct styles or layout compositions. By leveraging this knowledge, we can implement diversified sampling based on style or layout, offering users varied results. Lastly, our studies on image inpainting reveal that some seeds consistently generate ‘text artifacts’ instead of completing pixels, indicating that one could improve inpainting quality by using seeds that minimize these artifacts. Note that for all these applications, we only need to perform the seed analysis once per model, and our approach can be easily integrated into the inference process without adding any computational overhead, unlike most optimization-based approaches.
@@
  • We discovered that seeds encode highly discriminative information, enabling a classifier to easily predict the source seed from a pool of 1,024 possible seeds with 99.9% validation accuracy using only the generated image as input.
@@
  • Capitalizing on our insights, we propose downstream applications that enhance high-fidelity or diversified inference for text-to-image models, as well as improve generation quality by avoiding ‘text artifacts’ in text-based inpainting models.
@@ 2 Related Work
Impact of diffusion model inputs. The main sources of variation in images produced by pretrained text-to-image diffusion models [2, 3, 6, 31, 35, 36, 56] are the text prompt and the random seed that controls the initial noise. Consequently, carefully selecting these model inputs can enhance image generation and editing during inference without requiring additional model training or fine-tuning. Several studies [30, 47, 51, 55] have focused on understanding the impact of text embeddings on the generated image or leveraging these text embeddings for tuning-free image generation. For instance, Yu et al. [55] discovered that the CLIP [34] text embedding commonly used in T2I diffusion models contains diverse semantic directions that facilitate controllable image editing. Furthermore, recent works [14, 27, 28, 33] have shown that the initial noise can lead to certain image generation tendencies. In particular, Po-Yuan et al. [33] demonstrated that slight perturbations to the initial noise can substantially alter the generated samples of a diffusion model. However, the extent to which the initial noise affects various visual dimensions of the output image remains unclear. Therefore, we conduct an extensive analysis of the influence of random seeds on the generated image’s quality, human preference alignment, style, composition, and insertion of ‘text artifacts.’
Optimizing initial noise in diffusion models. Given the significant impact of the seed on images generated by T2I diffusion models, previous works [7, 14, 27, 28, 37] have aimed to optimize the initial noise to produce images that better align with the text prompt, reduce visual artifacts, or achieve a desired layout. For example, Mao et al. [28] found that certain patches of initial noise are more likely to denoise into specific concepts, enabling them to approach image editing by simply substituting regions of the initial noise without fine-tuning or disrupting the reverse diffusion process. While their work concentrates on a local analysis of the initial noise, our research provides a large-scale study of the random seeds that control the initial noise across a diverse set of text prompts.
@@ 3.1 What do seeds control in the reverse diffusion process?
We focus on latent diffusion models as described by Rombach et al. [36], although the same principles apply to pixel diffusion models. Theoretically, in the traditional multi-step reverse diffusion process, both the initial noisy latent variables and the noise used for reparameterization [18] at each timestep are sampled from a Gaussian distribution, introducing randomness. We visualize this process on the left side of Figure 2. At the implementation level, we confirmed that random seeds are used as inputs to compute these variables [50]. In a distilled one-step diffusion model, such as SDXL Turbo [38], the random seeds only determine the initial noisy latent, as there are no intermediate denoising steps.
In multi-step diffusion inference, seeds determine both the initial latent variables and the reparameterization noise at each timestep. To understand the separate impacts of the initial latent configuration and the reparameterization step on the generated images, we conducted a simple "seed swap" study shown on the right side of Figure 2 using the DDIM scheduler [45] with 40 inference steps. In our study, we first set the seed to $i$ and begin the reverse diffusion process. Then, at an intermediate timestep, we change the seed to $j$ and complete the image generation process. We explore using seeds 0 and 1 for both $i$ and $j$, as well as swapping the seed at early, mid, and late timesteps of the reverse diffusion process. Despite these variations, we found that the initial noisy latent significantly controls the generated content, while the random noise introduced at intermediate reparameterization steps has no visible impact on the generated images, as shown on the right side of Figure 2.
@@ 3.2 Data Generation
To conduct seed analysis at large scale, we employ three types of prompts for text-to-image (T2I) generation, as shown in Figure 3. First, to capture a broad spectrum of natural visual content, we sample 20,000 images from the commonly used MS-COCO 2017 train set [24] and generate dense captions using LLaVA 1.5 [25]. Second, we utilize 1,632 prompts from the PartiPrompts benchmark [56], which includes short and long general-purpose user prompts. Lastly, to enable more controlled scientific studies, we create synthetic prompts by combining 40 object categories with 22 modifiers, resulting in 880 unique combinations.
For each prompt in our dataset, we sample 1,024 seeds and generate images using two T2I models, SD 2.0 [36] and SDXL Turbo [38], for a large-scale seed analysis. This results in a total number of $22,512$ prompts ×1,024$\times$ 1,024× 1 , 024 seeds ×$\times 2$ models $=46,104,576$ images. Beyond text-to-image applications, we also curated 500 pairs of images and masks for diffusion inpainting models, where the mask typically covers an object in the original image. For the text prompts, we use "clear background" to simulate the object removal use case and the original object category for the object completion use case, where the details are discussed in Section 4.3.
@@ 3.3 How discriminative are seeds based on their generated images?
As an initial experiment, we examine whether seeds can be distinguished by their generated images. We train a 1,024-way classifier to predict the seed number used to produce a given image, employing 9,000 training, 1,000 validation, and 1,000 test images per seed. Remarkably, seeds are highly differentiable based on their images. After only six epochs, our classifier trained on images from SD 2.0 [36] achieved a test accuracy of 99.994%, and the classifier trained on images from SDXL Turbo [38] reached a test accuracy of 99.956%. However, it is unclear what makes seeds easily discernible, as the Grad-CAM [13, 41] visualization in Figure 4 is not easily interpretable. These findings suggest that seeds may encode unique visual features, prompting us to explore their impact across several interpretable dimensions.
3.4 Impact of Seeds on Interpretable Visual Dimensions
In Section 3.3, we observed that a classifier trained to predict the seed used to generate an image achieves over 99.9% accuracy in just a few epochs of training. However, it remains unclear what aspects of the generated images enable these seeds to be highly distinguishable. Therefore, we present an extensive empirical study on the influence of seed number on interpretable visual dimensions.
Image Quality and Human Preference Alignment. As mentioned in Section 3.2, we used 20,000 prompts from MS-COCO dense captions [24, 25]. For each prompt, we generated images using 1,024 seeds. To evaluate the image quality associated with each seed, we selected 10,000 prompts and their corresponding generated images, and then computed the FID score [15, 40] against 10,000 real MS-COCO images [24]. Surprisingly, we observed a significant difference in FID scores between the best and worst seeds. For instance, the ‘golden’ seed 469 for SD 2.0 achieved a low FID of 21.60, while the ‘inferior’ seed 696 scored 31.97—a disparity considered significant within the community. Additionally, we assess the seeds using HPS v2 [52], a new metric trained on large-scale human preference pairs to quantify human preferences for AI-generated images. For each seed, we sampled 1,000 prompts and their corresponding images to calculate HPS v2. As shown in Figure 5, the top and bottom three seeds according to FID and HPS v2 indeed reveal that the highest-rated seeds produce images that are more visually pleasing and aligned with human preferences.
@@ 3.4 Impact of Seeds on Interpretable Visual Dimensions
Next, we determine whether these seed rankings are generalizable across a different set of 10,000 prompts for FID and 1,000 prompts for HPS v2. In Figures 6 and 7, we plot the ranked seeds for FID and HPS v2 using images from SD 2.0 and SDXL Turbo. We compare scores from the first set of prompts ("Prompt Set 1" in blue) against scores from another set of prompts ("Prompt Set 2" in orange). We reveal a high degree of overlap between the seed patterns for quality and human preference, indicating that they are truly applicable across sets of prompts. This consistency underpins our proposed enhancements to inference strategies detailed in Sections 4.1 and 4.2.
@@
Image Style. Given the visual variations in images generated using different seeds, we investigate whether specific seeds consistently produce unique style patterns across various prompts. Drawing on established methods in image texture and style transfer [11, 12], we compute style representations by extracting the Gram matrix — which measures pairwise cosine similarity across channels — from a pretrained deep network [44] at multiple layers. Next, we reshape the Gram matrix into a single-column vector for each image and reduce its dimensionality to two using PCA and t-SNE [1, 49]. Now, for each image, we have a compact 2D vector that captures its style. For $N=1024$ seeds and $P$ prompts, this results in a feature dimension of $N\times(2\times P)$, combining the style representation across the generated images for each seed. We further reduce [1, 49] this aggregated style representation per seed from $N\times(2\times P)$ to $N\times 2$. Finally, a subset of seeds are visualized in 2D in Figure 8, providing a clear visual representation of style clustering at the seed level.
@@
In Figure 8, the positions within the embedding space correspond to the same seeds across various subplots. As depicted in the first row, certain seed groups consistently generate grayscale images irrespective of the prompts used. Similarly, the second row shows that some seeds tend to produce images with prominent sky regions, while others do not. Furthermore, in Figure 9, we observe that a select group of seeds consistently generate images with a ‘border’ effect near the edges, regardless of the text prompt. Collectively, these findings demonstrate that individual seeds exhibit distinct tendencies in style generation across varying prompts.
Figure 9: We observe that certain seeds produce a "border" around the image for SD 2.0. Often, these borders appear as horizontal bars at the top and bottom. Surprisingly, seed 0 occasionally generates a thick dark border on the left side of the image, while seed 50 sometimes adds a "photo frame."
Image Composition. Moving beyond style, we examine whether seeds create distinctive image compositions, such as consistent object locations and sizes. As described in Section 3.2, we generate images using 880 synthetic prompts consisting of 40 object categories paired with 22 modifiers, which includes adjectives and the empty string. For each image, we segment [8] the object and compute an image composition feature vector that contains the object’s centroid $(x,y)$ coordinates, size, and depth [53] relative to the image. On the left side of Figure 10, we visualize the distribution of the object mask’s centroid for the category "horse." Remarkably, the object’s position stays relatively the same despite slight prompt alterations. On the right side of Figure 10, we observe an analogous pattern in the object’s size and depth for the category "bowl." Overall, we observe that the location, size, and depth of generated objects are largely dependent on the specific seed used, consistent across the same object categories and irrespective of the text modifiers in the prompts.
Figure 10: We observe that seeds produce images with unique and consistent compositions for a given object category. Each data point represents a seed. For each seed, we combine image composition features from 22 prompts with slight variations like "a pale bowl" and "a round bowl." Then, we apply dimensionality reduction [1, 49] for visualization. Left: Distribution of object centroid $(x,y)$ coordinates. Right: Distribution of object depth and size relative to the image.

### v2

@@ Abstract
Recent advances in text-to-image (T2I) diffusion models have facilitated creative and photorealistic image synthesis. By varying the random seeds, we can generate many images for a fixed text prompt. Technically, the seed controls the initial noise and, in multi-step diffusion inference, the noise used for reparameterization at intermediate timesteps in the reverse diffusion process. However, the specific impact of the random seed on the generated images remains relatively unexplored. In this work, we conduct a large-scale scientific study into the impact of random seeds during diffusion inference. Remarkably, we reveal that the best ‘golden’ seed achieved an impressive FID of 21.60, compared to the worst ‘inferior’ seed’s FID of 31.97. Additionally, a classifier can predict the seed number used to generate an image with over 99.9% accuracy in just a few epochs, establishing that seeds are highly distinguishable based on generated images. Encouraged by these findings, we examined the influence of seeds on interpretable visual dimensions. We find that certain seeds consistently produce grayscale images, prominent sky regions, or image borders. Seeds also affect image composition, including object location, size, and depth. Moreover, by leveraging these ‘golden’ seeds, we demonstrate improved image generation such as high-fidelity inference and diversified sampling. Our investigation extends to inpainting tasks, where we uncover some seeds that tend to insert unwanted text artifacts. Overall, our extensive analyses highlight the importance of selecting good seeds and offer practical utility for image generation.
@@ 1 Introduction
Text-to-Image (T2I) diffusion models [2, 3, 6, 32, 36, 37, 57] have advanced image synthesis significantly, enabling the creation of photorealistic, high-resolution images. But, their training requires substantial compute, limiting such research to a few well-equipped labs. Despite these limitations, many studies have enhanced image generation during inference by feature re-weighting [44], gradient-based guidance [10, 43, 49], or fusion with multimodal LLMs [5, 55].
@@
In this work, we propose an inference technique to enhance image generation by exploring ‘secret seeds’ in the reverse diffusion process. Inspired by prior research [33], which revealed that well-chosen neural network initialization seeds can outperform poorly chosen ones in image classification, we investigate whether ‘golden’ or ‘inferior’ seeds similarly impact image quality in T2I diffusion inference. Surprisingly, using the pretrained T2I model Stable Diffusion (SD) 2.0 [37] across 1,024 seeds, we discovered that the best ‘golden’ seed achieved an FID [16, 41] of 21.60, whereas the worst ‘inferior’ seed only reached an FID of 31.97—a significant difference within the community. This finding sparked our curiosity to understand several scientific questions: What does the seed control in T2I diffusion inference? Why are random seeds so impactful? Can seeds be distinguished by the images they generate? Do they control interpretable image dimensions, and if so, how can this be leveraged to enhance image generation?
@@
To address our research questions, we first examined how random seeds control the initial noisy latent and the Gaussian noise during the reparameterization step of each intermediate timestep in the reverse latent diffusion process, as detailed in Sec. 3.1. We also developed a dataset using two T2I diffusion models: the conventional multi-step SD 2.0 [37] and the distilled one-step SDXL Turbo [39]. This dataset includes over 22,000 diverse text prompts and, using 1,024 unique fixed seeds for each combination of model and prompt, resulted in approximately 46 million images as discussed in Sec. 3.2. Our initial objective was to determine whether each random seed encodes unique characteristics identifiable in the generated images. To test this, we trained a 1,024-way classifier to predict the seed number used during diffusion inference from the generated images across diverse prompts. Remarkably, this classifier reached over 99.9% validation accuracy after just six epochs, a stark contrast to the random guessing chance of approximately 0.01%, establishing that seeds are highly distinguishable based on the generated images as shown in Sec. 3.3.
@@
Having confirmed seed distinguishability, we aim to understand if there are interpretable perceptual dimensions enabling this differentiation. Our next step involves designing a pipeline to extract style and layout representations, apply dimensionality reduction [1, 50] for visible clustering, and identify consistent patterns across seeds, regardless of the input prompts. For example, certain seeds consistently produce ‘grayscale’ images, others generate images with prominent white ‘sky’ regions at the top, and some seeds create image borders or insert ‘text’ during inpainting mode. In terms of image layout, various seeds consistently influence the main subject’s scale, location, and depth within images. The details on these findings are in Sec. 3.4.
@@
Building on these discoveries from our seed analysis, we propose several downstream applications to enhance image generation, as detailed in Sec. 4. First, by identifying ‘golden’ seeds across a variety of prompts, we can limit sampling to the top-K seeds for high-fidelity inference. This approach demonstrates superior quantitative performance, as measured by FID [41] and HPS v2 [53], compared to random sampling in the default implementation. Second, our findings indicate that certain seeds capture distinct styles or layout compositions. By leveraging this knowledge, we can implement diversified sampling based on style or layout, offering users varied results. Lastly, our studies on image inpainting reveal that some seeds consistently generate ‘text artifacts’ instead of completing pixels, indicating that one could improve inpainting quality by using seeds that minimize these artifacts. Note that for all these applications, we only need to perform the seed analysis once per model, and our approach can be easily integrated into the inference process without adding any computational overhead, unlike most optimization-based approaches.
@@
  • We discovered that seeds encode highly discriminative information, enabling a classifier to predict the source seed from 1,024 possible seeds with 99.9% validation accuracy using only the generated image as input.
@@
  • Capitalizing on our insights, we propose applications that enhance high-fidelity or diversified inference for T2I models, and improve generation quality by avoiding ‘text artifacts’ in text-based inpainting models.
@@ 2 Related Work
Impact of diffusion model inputs. The main sources of variation in images produced by pretrained text-to-image diffusion models [2, 3, 6, 32, 36, 37, 57] are the text prompt and the random seed that controls the initial noise. Consequently, carefully selecting these model inputs can enhance image generation and editing during inference without requiring additional model training or fine-tuning. Several studies [31, 48, 52, 56] have focused on understanding the impact of text embeddings on the generated image or leveraging these text embeddings for tuning-free image generation. Yu et al. [56] discovered that the CLIP [35] text embedding commonly used in T2I diffusion models contains diverse semantic directions that facilitate controllable image editing. Furthermore, recent works [15, 28, 29, 34] have shown that the initial noise can lead to certain image generation tendencies. Po-Yuan et al. [34] demonstrated that slight perturbations to the initial noise can greatly alter the generated samples of a diffusion model, and Grimal et al. studied the effect of seeds on text-image alignment [14]. However, the extent to which the initial noise affects various visual dimensions of the image remains unclear. Therefore, we conduct an extensive analysis of the influence of random seeds on the generated image’s quality, human preference alignment, style, composition, and insertion of ‘text artifacts.’
Optimizing initial noise in diffusion models. Given the significant impact of the seed on images from T2I diffusion models, previous works [7, 15, 28, 29, 38] have aimed to optimize the initial noise to produce images that better align with a text prompt, reduce visual artifacts, or achieve a desired layout. Mao et al. [29] found that certain patches of initial noise are more likely to denoise into specific concepts, enabling them to approach image editing by simply substituting regions of the initial noise without fine-tuning or disrupting the reverse diffusion process. While their work focuses on a local analysis of the initial noise, our research provides a large-scale study of the random seeds that control the initial noise across a diverse set of text prompts.
@@ 3.1 What do seeds control in the reverse diffusion process?
We focus on latent diffusion models as described by Rombach et al. [37], but the same principles apply to pixel diffusion models. Theoretically, in the traditional multi-step reverse diffusion process, both the initial noisy latent variables and the noise used for reparameterization [19] at each timestep are sampled from a Gaussian distribution, introducing randomness. We show this process on the left side of Fig. 2. At the implementation level, we confirmed that random seeds are inputs to compute these variables [51]. In a distilled one-step diffusion model, such as SDXL Turbo [39], the random seeds only determine the initial noisy latent, as there are no intermediate denoising steps.
In multi-step diffusion inference, seeds determine both the initial latent variables and the reparameterization noise at each timestep. To understand the separate impacts of the initial latent configuration and the reparameterization step on the generated images, we conducted a simple "seed swap" study shown on the right side of Fig. 2 using the DDIM scheduler [46] with 40 inference steps. In our study, we first set the seed to $i$ and begin the reverse diffusion process. Then, at an intermediate timestep, we change the seed to $j$ and complete the image generation process. We explore using seeds 0 and 1 for both $i$ and $j$, as well as swapping the seed at early, mid, and late timesteps of the reverse diffusion process. Despite these variations, we found that the initial noisy latent significantly controls the generated content, while the random noise introduced at intermediate reparameterization steps has no visible impact on the generated images, as shown on the right side of Fig. 2.
@@ 3.2 Data Generation
To conduct seed analysis at large scale, we employ three types of prompts for text-to-image (T2I) generation, as shown in Fig. 3. First, we capture a broad spectrum of natural visual content by sampling 20,000 images from the MS-COCO 2017 train set [25] and generate dense captions using LLaVA 1.5 [26]. Second, we utilize 1,632 prompts from the PartiPrompts benchmark [57], which includes short and long general-purpose user prompts. Lastly, to enable more controlled scientific studies, we create synthetic prompts by combining 40 object categories with 22 modifiers, resulting in 880 unique combinations.
For each prompt in our dataset, we sample 1,024 seeds ranging from 0 to 1,023 and generate images using two T2I models, SD 2.0 [37] and SDXL Turbo [39], for a large-scale seed analysis. Specifically, we assign the seed via torch.Generator(“cuda").manual_seed(seed) and use a DDIM scheduler [46] for SD 2.0. This results in a total number of $22,512$ prompts ×1,024$\times$ 1,024× 1 , 024 seeds ×$\times 2$ models $=46,104,576$ images. Beyond text-to-image applications, we curated 500 pairs of images and masks for diffusion inpainting models based on Open Images [22, 20], where the mask typically covers an object in the original image. For the text prompts, we use “clear background" to simulate the object removal case and the original object category for the object completion case, where the details are in Sec. 4.3.
@@ 3.3 How discriminative are seeds based on their generated images?
As an initial experiment, we examine whether seeds can be distinguished by their generated images. We train a 1,024-way classifier to predict the seed number used to produce an image, employing 9,000 training, 1,000 validation, and 1,000 test images per seed. Remarkably, seeds are highly differentiable based on their images. After only six epochs, our classifier trained on images from SD 2.0 [37] achieved a test accuracy of 99.99%, and the classifier trained on images from SDXL Turbo [39] reached a test accuracy of 99.96%. However, it is unclear what makes seeds easily discernible, as the Grad-CAM [13, 42] visualization in Fig. 4 is not clearly interpretable. These findings suggest that seeds may encode unique visual features; thus, we explore their impact across several interpretable dimensions.
@@
3.4 Impact of Seeds on Interpretable Dimensions
In Sec. 3.3, we observed that a classifier trained to predict the seed used to generate an image achieves over 99.9% accuracy in just a few epochs of training. However, it remains unclear what aspects of the generated images enable these seeds to be highly distinguishable. Therefore, we present an extensive empirical study on the influence of seed number on interpretable visual dimensions.
@@ 3.4 Impact of Seeds on Interpretable Dimensions
Image Quality and Human Preference Alignment. As mentioned in Sec. 3.2, we used 20,000 prompts from MS-COCO dense captions [25, 26]. For each prompt, we generated images using 1,024 seeds. To evaluate the image quality associated with each seed, we selected 10,000 prompts and their corresponding generated images, and then computed the FID score [16, 41] against 10,000 real MS-COCO images [25]. Surprisingly, we observed a significant difference in FID scores between the best and worst seeds. For instance, the ‘golden’ seed 469 for SD 2.0 achieved a low FID of 21.60, while the ‘inferior’ seed 696 scored 31.97—a disparity considered significant within the community. Additionally, we assess the seeds using HPS v2 [53], a new metric trained on large-scale human preference pairs to quantify human preferences for AI-generated images. For each seed, we sampled 1,000 prompts and their corresponding images to calculate HPS v2. As shown in Fig. 5, the top and bottom three seeds according to FID and HPS v2 indeed reveal that the highest-rated seeds produce images that are more visually pleasing and aligned with human preferences.
Next, we determine whether these seed rankings are generalizable across a different set of 10,000 prompts for FID and 1,000 prompts for HPS v2. In Fig. 6 and 7, we plot the ranked seeds for FID and HPS v2 using images from SD 2.0 and SDXL Turbo. We compare scores from the first set of prompts (“Prompt Set 1") against scores from another set of prompts (“Prompt Set 2"). We reveal a high degree of overlap between the seed patterns for quality and human preference. This consistency underpins our proposed enhancements to inference strategies in Sec. 4.1 and 4.2.
Image Style. Given the visual variations in images generated using different seeds, we investigate whether specific seeds consistently produce unique style patterns across prompts. Drawing on established methods in image texture and style transfer [11, 12], we compute style representations by extracting the Gram matrix — which measures pairwise cosine similarity across channels — from a pretrained deep network [45] at multiple layers. Next, we reshape the Gram matrix into a single-column vector for each image and reduce its dimensionality to two using PCA and t-SNE [1, 50]. Now, for each image, we have a compact 2D vector that captures its style. Using $N=1024$ seeds and $P$ prompts results in a feature dimension of $N\times(2\times P)$, combining the style representation across the generated images for each seed. We further reduce [1, 50] this aggregated style representation per seed from $N\times(2\times P)$ to $N\times 2$. Finally, a subset of seeds are visualized in Fig. 8, providing a clear visual representation of style clustering at the seed level.
@@
In Fig. 8, the positions within the embedding space correspond to the same seeds across various subplots. As depicted in the first row, certain seed groups consistently generate grayscale images irrespective of the prompt. Similarly, the second row shows that some seeds tend to produce images with prominent sky regions, while others do not. Furthermore, in Fig. 9, we observe that a select group of seeds consistently generate images with a ‘border’ effect near the edges, regardless of the text prompt. Collectively, these findings demonstrate that individual seeds exhibit distinct tendencies in style generation across varying prompts.
Figure 9: Certain seeds produce a “border" around images for SD 2.0. Often, these borders appear as horizontal bars at the top and bottom. Surprisingly, seed 0 occasionally generates a thick dark border on the left side of the image, while seed 50 sometimes adds a “photo frame."
Image Composition. Moving beyond style, we examine whether seeds create distinctive image compositions, such as consistent object locations and sizes. As described in Sec. 3.2, we generate images using 880 synthetic prompts consisting of 40 object categories paired with 22 modifiers, which includes adjectives and the empty string. For each image, we segment [8] the object and compute an image composition feature vector that contains the object’s centroid $(x,y)$ coordinates, size, and depth [54] relative to the image. On the left side of Fig. 10, we visualize the distribution of the object mask’s centroid for the category “horse." Remarkably, the object’s position stays relatively the same despite slight prompt alterations. On the right side of Fig. 10, we observe an analogous pattern in the object’s size and depth for the category “bowl." Overall, we observe that the location, size, and depth of objects are largely dependent on the specific seed used, consistent across the same object categories and irrespective of the text modifiers in the prompts.
Figure 10: We observe that seeds produce images with unique and consistent compositions for a given object category. Each data point represents a seed. For each seed, we combine image composition features from 22 prompts with slight variations like “a pale bowl" and “a round bowl." Then, we apply dimensionality reduction [1, 50] for visualization. Left: Distribution of object centroid $(x,y)$ coordinates. Right: Distribution of object depth and size relative to the image.